A rectangular floor has an area of 24 square meters. The length of the floor is 2 meters less than twice the width. What is the length and width of the floor
Answer by amarjeeth123(569)

Let the width be x.
Then the length is (2x-2).
Area=24 square metres
Area=length*width
(2x-2)x=24
x(x-1)=12
x^2-x-12=0
x^2-4x+3x-12=0
x(x-4)+3(x-4)=0
(x+3)(x-4)=0
x=4
The width of the floor is 4 metres.
The length of the floor is 6 metres.